# Development of a conditional diffusion model to predict process parameters and microstructures of dendrite crystals of matrix resin based on mechanical properties

### v1

@@ Abstract
In this study, we develop a conditional diffusion model that proposes the optimal process parameters, such as processing temperature, and predicts the microstructure for the desired mechanical properties, such as the elastic constants of the matrix resin contained in carbon fiber reinforced thermoplastics (CFRTPs). In CFRTPs, not only the carbon fibers but also the matrix resin contribute to the macroscopic mechanical properties. Matrix resins contain a mixture of dendrites, which are crystalline phases, and amorphous phases even after crystal growth is complete, and it is important to consider the microstructures consisting of the crystalline structure and the remaining amorphous phase to achieve the desired mechanical properties. Typically, the temperature during forming affects the microstructures, which in turn affect the macroscopic mechanical properties. The training data for the conditional diffusion model in this study are the crystallization temperatures, microstructures and the elasticity matrix. The elasticity matrix is normalized and introduced into the model as a condition. The trained diffusion model can propose not only the processing temperature but also the microstructure when Young’s modulus and Poisson’s ratio are given. The capability of our conditional diffusion model to represent complex dendrites is also noteworthy.
keywords:
conditional diffusion model , extended finite element method , phase-field method , dendrite , matrix resin
[…]
[inst1]organization=Graduate School of Science and Technology, Keio University,addressline=3-14-1, city=Hiyoshi, Kohoku-ku, Yokohama, postcode=223-8522, state=Kanagawa, country=Japan
[…]
[inst2]organization=Department of Aeronautics and Astronautics, The University of Tokyo,addressline=7-3-1, city=Hongo, Bunkyo-ku, postcode=113-8656, state=Tokyo, country=Japan
[…]
[inst3]organization=National Institute of Advanced Industrial Science and Technology,addressline=1-1-1, city=Umezono, Tsukuba, postcode=305-8568, state=Ibaraki, country=Japan
[…]
[inst4]organization=Department of Mechanical Engineering, Keio University,addressline=3-14-1, city=Hiyoshi, Kohoku-ku, Yokohama, postcode=223-8522, state=Kanagawa, country=Japan
@@ 1 Introduction
Carbon fiber reinforced thermosetting plastics (CFRTSs) and carbon fiber reinforced thermoplastics (CFRTPs) have excellent light weight and strength [1]. These materials are increasingly applied to structural components of aircraft and automobiles [2, 3]. However, aircraft and automobiles have different requirements. Therefore, it is important to develop materials with the desired properties for the application. CFRTSs have the property of stiffening upon heating and not returning to its original state, whereas CFRTPs have the property of stiffening upon cooling and softening again upon heating. CFRTSs have been applied to many structures, but the material and molding costs are high. In addition, it is difficult to recycle CFRTSs once stiffened, which poses a disposal problem [4, 5]. On the other hand, CFRTPs have been attracting attention as a sustainable material due to its low molding cost and recyclability [6]. In CFRTPs, the mechanical properties depend not only on the carbon fibers, but also on the matrix resin used as the base material. The thermoplastic resins used in CFRTPs are crystalline resins, such as polyphenylene sulfide (PPS) and polyether ether ketone (PEEK), which have significant rigidity [7, 8]. Crystalline and amorphous phases are mixed in matrix resins, and it is important to consider the microstructures consisting of the crystalline structure and the remaining amorphous phase to achieve the desired mechanical properties. As crystallization progresses, matrix resins become stronger and stiffer. This increases the overall strength and stiffness of the composite. On the other hand, the amorphous phase contributes to the ductility of the composite. Therefore, it is important to consider the balance of crystalline and amorphous phase and their arrangement.
In this study, we focus on PPS, a thermoplastic resin with low costs. PPS has a mixture of crystalline and amorphous phases even when the crystals are fully grown. It produces dendritic crystals called dendrites during solidification [9]. Process parameters such as temperature during forming and cooling rate affect the microstructure, which in turn affects the macroscopic mechanical properties [10, 11, 12, 13, 14, 15, 16]. Thus, a method of predicting the mechanical properties of resins is required. To improve the mechanical properties of composites such as CFRTP and CFRTS, various experiments have been tried [17, 18, 19, 20, 21, 22, 23]. It is important to investigate the properties of materials through experimentation. On the other hand, it is very expensive to try different conditions in the experiment. To solve this problem and to develop materials efficiently, simulation has been introduced.
@@
Higuchi et al. [24] and Takashima et al. [25] developed a multiphysics analysis method for crystalline thermoplastic resins that links the forming conditions, the microstructure and the macroscopic mechanical properties. Crystallization analysis was performed by using the phase-field method [26, 27, 28, 29]. The homogenization analysis [30] of PPS, a crystalline thermoplastic resin, was conducted by using the extended finite element method (XFEM) [31, 32, 33, 34, 35, 36]. PPS was characterized to identify the parameters necessary for crystallization analysis. The simulation results were validated by comparing them with experimental data. The details of their studies [24, 25] are as follows. They measured properties such as the glass transition temperature, crystallization temperature and melting point by differential scanning calorimetry (DSC). Moreover, they examined the relationship between the processing temperature and the mechanical properties by tensile tests. The phase-field model they developed reproduced various types of nucleation and polycrystal growth. Conventional methods for dendrite simulation divide the analysis domain into cells and discretely assign states to them, e.g., the cellular automata [37, 38, 39, 40] and Monte Carlo [41, 42] methods. The other method represents the interface by connecting discrete points such as the front-tracking method [43, 44, 45]. Higuchi et al. [24] used the phase-field method to describe the crystal growth of dendrites in thermoplastic resins, in terms of the interface movement between two different phases. Then, they performed the homogenization analysis of the obtained microstructures using XFEM to determine the mechanical properties such as Young’s modulus. By comparing the analytical results with the experimental results, it is shown that their proposed method has a certain validity. Their method is limited to forward analysis, and it is not possible to predict the processing temperature or microstructure from the mechanical properties. In addition, machine learning is sometimes used for efficient material development because it takes a long time to try many patterns in a simulation.
In the field of mechanics of materials, machine learning has achieved significant results in the prediction of physical properties [46, 47, 48, 49, 50, 51, 52]. Jadhav et al. [46] proposed a StressD framework for predicting von Mises stress fields of structures based on the denoising diffusion model. They applied the StressD framework to four datasets. The datasets [53, 54] are coarse mesh, fine mesh, sparse and hyperelastic datasets. The coarse mesh dataset has cantilever structures composed of homogeneous and isotropic linear elastic materials. The fine mesh dataset has 50,000 data points sampled from the coarse mesh dataset. The sparse dataset has data randomly sampled from $5\%$ of the stress field in the material domain. The hyper elastic dataset has distinctive data based on a two-dimensional (2D) hyperelastic problem. The StressD framework has two models: the denoising diffusion model and auxiliary network. The denoising diffusion model generates a normalized stress map on the basis of a given geometry and other conditions. The auxiliary network determines the scaling information required to un-normalize the generated stress map. Compared with the conventional method of finite element analysis (FEA), the computational cost of the StressD framework is significantly reduced and the von Mises stress field is accurately predicted. In this way, machine learning contributes to the reduction of computational cost by using it for the prediction of physical properties.
In material development, machine learning has also contributed markedly [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. Kumar et al. [55] developed a machine learning strategy for the inverse design of spinodoid topology. This strategy enables the creation of metamaterials with specific mechanical properties. Bastek et al. [56] developed a method for the inverse design of nonlinear mechanical metamaterials using a video denoising diffusion model. The video denoising diffusion model is trained on full-field data of periodic stochastic cellular structures. They showed that the model can predict and tune the nonlinear deformation and stress response under compression in the large-strain regime. Veasna et al. [57] developed a Pareto-based multiobjective machine learning method for efficiently identifying parameters of crystal plasticity models. The model parameters are identified for two steel alloys, the dual-phase (DP) steel DP780 and the stainless steel (SS) SS316L. Their method greatly reduces the number of expensive black-box function runs and allows the parameters to be accurately specified. Hiraide et al. [58] developed a framework for forward analysis to predict Young’s modulus from the phase-separated structure of polymer alloys and for inverse analysis to predict the structure from Young’s modulus. Forward analysis uses convolutional neural networks (CNNs) [66, 67]. For inverse analysis, random search is applied to the combined generative adversarial network (GAN) [68, 69] and CNN model. Moreover, Hiraide et al. [59] proposed a framework for designing material structures based on macroscopic properties. They use the results of the analysis of the two-dimensional phase separation structure of diblock copolymer melts as structural data. The stress data are obtained by FEA. The framework consists of a deep learning model that generates structures and a model that predicts the physical properties of the structures. It generates structures with desired physical properties using random search.
The denoising diffusion model, one of the machine learning models, is increasingly used in the design of materials [64]. Lee and Yun [64] proposed a framework for reconstructing anisotropic microstructures from 2D micrographs. They use the dataset from experimentally obtained 2D micrographs. The materials are a lithium-ion battery separator [70], carbon fiber composites [71] and lithium-ion nickel manganese cobalt (NMC) cathode samples [72]. To train conditional diffusion-based generative models (DGMs) for 2D-to-three-dimensional (3D) reconstruction, 2D micrographs from three orthogonal planes of each material are randomly sampled from the open-access tomographic data [70, 71, 72]. In this framework, 3D microstructures are generated from multiple 2D datasets using conditional DGMs. They validate the 3D reconstructed anisotropic microstructure samples by evaluating spatial correlation functions and physical material behavior.
@@
Thus, machine learning has been used to predict physical properties and develop materials. Methods of predicting mechanical properties from microstructures and microstructures from mechanical properties have been developed. However, there is no scheme for proposing process parameters and material microstructures related to how materials should be made. If process parameters such as processing temperature and cooling rate are proposed to produce materials with the desired mechanical property, it will be possible to develop materials efficiently by eliminating the need for trial-and-error experiments using different parameters.
@@
The purpose of this study is to develop a conditional diffusion model that proposes the optimal processing temperature and predicts the microstructure on the basis of the desired Young’s modulus and Poisson’s ratio. As an example, we analyzed PPS, a thermoplastic resin that is the base material of CFRTP, assuming composite applications.
@@ 2.1.3 Generation of crystal microstructure by the phase-field method
To generate the microstructure of matrix resins for machine learning, we use the phase-field method. The size of a nucleus is equal to one grid for simplicity. The analysis is performed at three crystallization temperatures: 160   C, 180   C and 200   C. Table 1 shows the conditions used in the analysis by the phase-field method. The condition parameters in the phase-field method are obtained from the references [29, 76].
@@ 2.2.1 XFEM
The elasticity matrix $\boldsymbol{D}$ used as the condition of the diffusion model is obtained by homogenization analysis using XFEM for matrix resins generated by the phase-field method. At the interface between the crystalline and amorphous phases, the strain is discontinuous and the local modification of the interpolation function allows accurate approximation. The nodes for XFEM are generated on the basis of the differentially discretized grid for the phase-field method and the triangular elements are generated by dividing the square grid into two parts. In addition, subelements are introduced for elements containing interfaces to integrate discontinuous functions with high accuracy. In XFEM, the displacement is expressed as
@@ 2.3 Conditional diffusion model
where $p$ is the stochastic process of the reverse process, $p_{\theta}$ is the stochastic process of the reverse process using the neural network and $\boldsymbol{\mu}_{\theta}$ is parameterized as a neural network. As shown in Fig. 3, the training data are the microstructures of matrix resins generated by the phase-field method, the processing temperatures specified in the phase-field analysis and the elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM for the microstructure generated by the phase-field analysis. For the training data of processing temperature, we generate image patterns of processing temperature (IPPT). Note that IPPT does not mean the distribution of temperatures, but gives a pattern for temperature conditions.
@@
Figure 4: Structure of ST-UNet [87]
The training data are the microstructures of matrix resins, IPPTs and the elasticity matrix $\boldsymbol{D}$. The microstructures are generated by the phase-field method, IPPT means the temperature specified in the phase-field analysis and the elasticity matrix $\boldsymbol{D}$ is obtained by homogenization analysis using XFEM for the microstructures generated by the phase-field method. We employ images of microstructures upon crystal growth completion. Images of the microstructures of dendrites are compressed for use in training machine learning. The mean values of $5{\ \mathrm{pixel}}\times 5{\ \mathrm{pixel}}$ regions are calculated for each $320{\ \mathrm{pixel}}\times 320{\ \mathrm{pixel}}$ analysis grid to be compressed to $64{\ \mathrm{pixels}}\times 64{\ \mathrm{pixels}}$. In addition, continuous phase-field variables are binarized during compression. Moreover, we perform data augmentation on the microstructures generated by the phase-field method. The $180^{\circ}$ rotation, inversion with respect to the $x_{1}$ axis, inversion with respect to the $x_{1}$ axis and $180^{\circ}$ rotation are conducted for the same elasticity matrix $\boldsymbol{D}$ value.
@@ 3.2 Validation of proposed processing temperatures and microstructures
Figs. 8 and 9 show the relationship between the crystallization temperature specified in the phase-field analysis and the elasticity matrix $\boldsymbol{D}$ obtained by XFEM for the microstructures generated by the phase-field method for (a) training and (b) test data, respectively. The violin plots in Fig. 8 show the distribution of the data with the horizontal axis for the crystallization temperature and the vertical axis for the value of each component of the elasticity matrix $\boldsymbol{D}$. Each violin plot shows that there is a large amount of data in the bulging part of the plot and a large scatter of data in the vertically long part of the plot. The black rectangle in the plot represents the interquartile range of the box plot, and the white dot at the center represents the median value. Fig. 9 shows scatter plots of $D_{1111}$ on the horizontal axis and $D_{2222}$ on the vertical axis, with different colors indicating the crystallization temperature. These plots show that the higher the crystallization temperature, the larger the value of each component of the elasticity matrix $\boldsymbol{D}$.
@@
Firstly, we confirmed the trained model by using the condition of the test data. Figs. 1014 show the results. Fig. 10 shows the generated images of microstructures and Fig. 11 shows the generated images that indicate the processing temperature. Fig. 12 shows the relationship between the proposed temperature and the elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM again for the generated microstructure. Fig. 13 is a scatter plot with $D_{1111}$ on the horizontal axis and $D_{2222}$ on the vertical axis, plotted in different colors for different crystallization temperature. From Figs. 12 and 13, the relationship between crystallization temperature and elasticity matrix $\boldsymbol{D}$ is generally appropriate because the tendencies of the training data and test data are similar. Fig. 14 compares the values of each component of the elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM of the generated microstructure again with the condition given for generation. The red line indicates the ideal situation that the elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM for the generated microstructure is equal to the elasticity matrix $\boldsymbol{D}$ given as the condition, i.e., when $D_{\text{XFEM}}=D_{\text{input}}$. Where $D_{\text{XFEM}}$ is the element of the elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM of the generated microstructure again, $D_{\text{input}}$ is the element of the elasticity matrix $\boldsymbol{D}$ given for generation. From the correlation coefficients and each plot, the elasticity matrix $\boldsymbol{D}$ obtained by the homogenization analysis using XFEM shows good agreement with the condition given for generation. The correlation coefficients indicate a strong positive correlation, which is considered that the predicted microstructure is appropriate for the given condition.

### v2

# Conditional diffusion model for inverse prediction of process parameters and dendritic microstructures from mechanical properties

@@ Abstract
In this study, we develop a conditional diffusion model that proposes the optimal process parameters and predicts the microstructure for the desired mechanical properties. In materials development, it is costly to try many samples with different parameters in experiments and numerical simulations. The use of data-driven inverse design method can reduce the cost of materials development. This study develops an inverse analysis model that predicts process parameters and microstructures. This method can be used for any material, but in this study it is applied to polymeric material, which is the matrix resin of carbon fiber reinforced thermoplastics as an example. Matrix resins contain a mixture of dendrites, which are crystalline phases, and amorphous phases even after crystal growth is complete, and it is important to consider the microstructures consisting of the crystalline structure and the remaining amorphous phase to achieve the desired mechanical properties. Typically, the temperature during forming affects the microstructures, which in turn affect the macroscopic mechanical properties. The trained diffusion model can propose not only the processing temperature but also the microstructure when Young’s modulus and Poisson’s ratio are given. The capability of our conditional diffusion model to represent complex dendrites is also noteworthy.
@@ 1 Introduction
In materials development, experiments require trial and error with multiple parameters, which is costly. To solve this problem, numerical simulations have been studied in various fields. In order to accelerate the computation by numerical simulation, hardware performance-based approaches such as parallel computing using graphics processing units parallelism[1] and software-based approaches using machine learning[2] have been used. For example, machine learning is used to predict mechanical properties [3, 4, 5, 6, 7]. These precious numerical simulations with machine learning are limited to forward analyses. Since forward analysis cannot infer the causes from effects, a data-driven approach to solve the inverse problem is required for more efficient material development. Hashemi et al. [8] presented a supervised machine learning based computational methodology for the design of particulate multifunctional composite materials with desired thermal conductivity, using the design of particulate composites with liquid metal elastomer as a case study. The method consists of three phases: data generation in the first phase, discovery of complex relationship between the structure and properties using appropriate machine learning algorithms in the second phase, and inference of direct structure-property relationships and generation and visualization of candidate microstructures by inverse design framework in the third phase. This new supervised machine learning approach accelerates the prediction of the thermal conductivity of particle composites and enables the design of composites with desirable properties. Lee et al. [9] aimed to provide a solution for inverse design even when data quality does not meet high standards, and presented a new design strategy employing two independent approaches: a metaheuristic-assisted inverse reading of conventional forward machine learning models and an atypical inverse machine learning model based on a modified variational autoencoder. They pinpointed several novel thermo-mechanically controlled processed steel alloy candidates, which were validated by a rule-based thermodynamic calculation tool. Bastek et al. [10] developed a method for the inverse design of nonlinear mechanical metamaterials using a video denoising diffusion model. The video denoising diffusion model is trained on full-field data of periodic stochastic cellular structures. They showed that the model can predict and tune the nonlinear deformation and stress response under compression in the large-strain regime. Hiraide et al. [11] developed a framework for forward analysis to predict Young’s modulus from the phase-separated structure of polymer alloys and for inverse analysis to predict the structure from Young’s modulus. Forward analysis uses convolutional neural networks (CNNs). For inverse analysis, random search is applied to the combined generative adversarial network (GAN) and CNN model. Moreover, Hiraide et al. [12] proposed a framework for designing material structures based on macroscopic properties. They use the results of the analysis of the two-dimensional phase separation structure of diblock copolymer melts as structural data. The stress data are obtained by the finite element analysis. The framework consists of a deep learning model that generates structures and a model that predicts the physical properties of the structures. It generates structures with desired physical properties using random search. Vlassis and Sun [13] presented a denoising diffusion algorithm to discover microstructures with nonlinear fine-tuned properties using the open-source mechanical MNIST data set provided in Lejeune[14]. They used a CNN architecture and a denoising diffusion algorithm, where the CNN architecture predicts the hyperelastic energy functional behavior under uniaxial extension, and the denoising diffusion algorithm generates targeted microstructures with desired constitutive responses. In this way, several data-driven approaches for inverse analysis were proposed for materials development.
@@
Carbon fiber reinforced thermosetting plastics (CFRTSs) and carbon fiber reinforced thermoplastics (CFRTPs) have excellent light weight and strength [15]. These materials are increasingly applied to structural components of aircraft and automobiles [16, 17]. CFRTSs have the property of stiffening upon heating and not returning to its original state, whereas CFRTPs have the property of stiffening upon cooling and softening again upon heating. CFRTSs have been applied to many structures, but the material and molding costs are high. In addition, it is difficult to recycle CFRTSs once stiffened, which poses a disposal problem [18, 19]. On the other hand, CFRTPs have been attracting attention as a sustainable material due to its low molding cost and recyclability. In CFRTPs, the mechanical properties depend not only on the carbon fibers, but also on the thermoplastic resin used as the base material. The thermoplastic resins used in CFRTPs are crystalline resins, such as polyphenylene sulfide (PPS) and polyether ether ketone (PEEK), which have significant rigidity [20, 21]. Crystalline and amorphous phases are mixed in thermoplastic resins, and it is important to consider the microstructures consisting of the crystalline structure and the remaining amorphous phase to achieve the desired mechanical properties. As crystallization progresses, thermoplastic resins become stronger and stiffer. This increases the overall strength and stiffness of the composite. On the other hand, the amorphous phase contributes to the ductility of the composite. Therefore, it is important to consider the balance of crystalline and amorphous phase and their arrangement. In this study, we focus on PPS, a thermoplastic resin with low costs. PPS has a mixture of crystalline and amorphous phases even when the crystals are fully grown. It produces dendritic crystals called dendrites during solidification [22]. Process parameters such as temperature during forming and cooling rate affect the microstructure, which in turn affects the macroscopic mechanical properties [23, 24, 25, 26, 27, 28]. Thus, a method of predicting the mechanical properties of resins is required.
@@
Higuchi et al. [29] and Takashima et al. [30] developed a multiphysics analysis method for crystalline thermoplastic resins that links the forming conditions, the microstructure and the macroscopic mechanical properties. Crystallization analysis was performed by using the phase-field method [31, 32, 33]. The homogenization analysis [34] of PPS, a crystalline thermoplastic resin, was conducted by using the extended finite element method (XFEM) [35, 36, 37]. PPS was characterized to identify the parameters necessary for crystallization analysis. The simulation results were validated by comparing them with experimental data. The details of their studies [29, 30] are as follows. They measured properties such as the glass transition temperature, crystallization temperature and melting point by differential scanning calorimetry (DSC). Moreover, they examined the relationship between the processing temperature and the mechanical properties by tensile tests. The phase-field model they developed reproduced various types of nucleation and polycrystal growth. Conventional methods for dendrite simulation divide the analysis domain into cells and discretely assign states to them, e.g., the cellular automata [38, 39, 40] and Monte Carlo [41, 42] methods. The other method represents the interface by connecting discrete points such as the front-tracking method [43, 44]. Higuchi et al. [29] used the phase-field method to describe the crystal growth of dendrites in thermoplastic resins, in terms of the interface movement between two different phases. Then, they performed the homogenization analysis of the obtained microstructures using XFEM to determine the mechanical properties such as Young’s modulus. By comparing the analytical results with the experimental results, it is shown that their proposed method has a certain validity. Their method is limited to forward analysis, and it is not possible to predict the processing temperature or microstructure from the mechanical properties.
@@
Thus, methods of predicting mechanical properties from microstructures and microstructures from mechanical properties have been developed. However, there is no scheme for proposing process parameters and material microstructures related to how materials should be made. If process parameters such as processing temperature and cooling rate are proposed to produce materials with the desired mechanical property, it will be possible to develop materials efficiently by eliminating the need for trial-and-error experiments using different parameters.
@@
The purpose of this study is to develop a conditional diffusion model that proposes the optimal process parameters and predicts the microstructure on the basis of the desired mechanical properties. As shown in Fig. 1, the inverse analysis model is applied to thermoplastic resins in this study to suggest the process temperature and predict the microstructure for the desired Young’s modulus and Poisson’s ratio.
@@ 2.1.3 Generation of crystal microstructure by the phase-field method
To generate the microstructure of thermoplastic resins for machine learning, we use the phase-field method. The size of a nucleus is equal to one grid for simplicity. The analysis is performed at three crystallization temperatures: 160   C, 180   C and 200   C. Table 1 shows the conditions used in the analysis by the phase-field method. The condition parameters in the phase-field method are obtained from the references [33, 48].
@@ 2.2.1 XFEM
The elasticity matrix $\boldsymbol{D}$ used as the condition of the diffusion model is obtained by homogenization analysis using XFEM for thermoplastic resins generated by the phase-field method. At the interface between the crystalline and amorphous phases, the strain is discontinuous and the local modification of the interpolation function allows accurate approximation. The nodes for XFEM are generated on the basis of the differentially discretized grid for the phase-field method and the triangular elements are generated by dividing the square grid into two parts. In addition, subelements are introduced for elements containing interfaces to integrate discontinuous functions with high accuracy. In XFEM, the displacement is expressed as
@@ 2.3 Conditional diffusion model
where $p$ is the stochastic process of the reverse process, $p_{\theta}$ is the stochastic process of the reverse process using the neural network and $\boldsymbol{\mu}_{\theta}$ is parameterized as a neural network. As shown in Fig. 4, the training data are the microstructures of thermoplastic resins generated by the phase-field method, the processing temperatures specified in the phase-field analysis and the elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM for the microstructure generated by the phase-field analysis. For the training data of processing temperature, we generate image patterns of processing temperature (IPPT). Note that IPPT does not mean the distribution of temperatures, but gives a pattern for temperature conditions.
@@
The training data are the microstructures of thermoplastic resins, IPPTs and the elasticity matrix $\boldsymbol{D}$. The microstructures are generated by the phase-field method, IPPT means the temperature specified in the phase-field analysis and the elasticity matrix $\boldsymbol{D}$ is obtained by homogenization analysis using XFEM for the microstructures generated by the phase-field method. We employ images of microstructures upon crystal growth completion. Images of the microstructures of dendrites are compressed for use in training machine learning. The mean values of $5{\ \mathrm{pixel}}\times 5{\ \mathrm{pixel}}$ regions are calculated for each $320{\ \mathrm{pixel}}\times 320{\ \mathrm{pixel}}$ analysis grid to be compressed to $64{\ \mathrm{pixels}}\times 64{\ \mathrm{pixels}}$. In addition, continuous phase-field variables are binarized during compression. Moreover, we perform data augmentation on the microstructures generated by the phase-field method. The $180^{\circ}$ rotation, inversion with respect to the $x_{1}$ axis, inversion with respect to the $x_{1}$ axis and $180^{\circ}$ rotation are conducted for the same elasticity matrix $\boldsymbol{D}$ value.
@@ 3.2 Validation of proposed processing temperatures and microstructures
Figs. 9 and 10 show the relationship between the crystallization temperature specified in the phase-field analysis and the elasticity matrix $\boldsymbol{D}$ obtained by XFEM for the microstructures generated by the phase-field method for (a) training and (b) test data, respectively. The violin plots in Fig. 9 show the distribution of the data with the horizontal axis for the crystallization temperature and the vertical axis for the value of each component of the elasticity matrix $\boldsymbol{D}$. Each violin plot shows that there is a large amount of data in the bulging part of the plot and a large scatter of data in the vertically long part of the plot. The black rectangle in the plot represents the interquartile range of the box plot, and the white dot at the center represents the median value. Fig. 10 shows scatter plots of $D_{1111}$ on the horizontal axis and $D_{2222}$ on the vertical axis, with different colors indicating the crystallization temperature. These plots show that the higher the crystallization temperature, the larger the value of each component of the elasticity matrix $\boldsymbol{D}$.
Firstly, we confirmed the trained model by using the condition of the test data. Figs. 1115 show the results. Fig. 11 shows the generated images of microstructures and Fig. 12 shows the generated images that indicate the processing temperature. Fig. 13 shows the relationship between the proposed temperature and the elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM again for the generated microstructure. Fig. 14 is a scatter plot with $D_{1111}$ on the horizontal axis and $D_{2222}$ on the vertical axis, plotted in different colors for different crystallization temperature. From Figs. 13 and 14, the relationship between crystallization temperature and elasticity matrix $\boldsymbol{D}$ is generally appropriate because the tendencies of the training data and test data are similar. Fig. 15 compares the values of each component of the elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM of the generated microstructure again with the condition given for generation. The red line indicates the ideal situation that the elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM for the generated microstructure is equal to the elasticity matrix $\boldsymbol{D}$ given as the condition, i.e., when $D_{\text{XFEM}}=D_{\text{input}}$. Where $D_{\text{XFEM}}$ is the element of the elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM of the generated microstructure again, $D_{\text{input}}$ is the element of the elasticity matrix $\boldsymbol{D}$ given for generation. From the correlation coefficients and each plot, the elasticity matrix $\boldsymbol{D}$ obtained by the homogenization analysis using XFEM shows good agreement with the condition given for generation. The correlation coefficients indicates a strong positive correlation, which is considered that the predicted microstructure is appropriate for the given condition.